OpenStudy (coolio100):

The two equal sides of an isosceles right triangle are of length z. Find a formula for the length of the hypotenuse

OpenStudy (welshfella):

Use the Pythagoras theorem

OpenStudy (welshfella):



OpenStudy (coolio100):

@welshfella so would it be c^2=z^2+z^2

OpenStudy (welshfella):

yes if c is the hypotenuse that is correct Now you have c^2 . Now find c.

OpenStudy (sshayer):

c^2=2 z^2 \[c=z \sqrt{2}\]